what is the formula for this sequence 0, 2, 5, 9, 14, 20 .......
Answer by mathstutor458(57)

0, 2, 5, 9, 14, 20 .......
0+2=2
2+3=5
5+4=9
9+5=14
14+6=20
20+7=27
therefore 2,3,4,5,6,7,----is adding to the numbers